a grain retailer puts a 1000 kg lead block under the platform of this scale. by how much does this increase the reading of the scale when 50kg sack of rice is on top of its platform, if the distance between their centers are .03 meter apart


1
Expert's answer
2022-01-18T09:36:11-0500

Solution


Mass of lead block m=1000kg

Mass of sack rice m'=50kg

Distance between them d=0.3m

We know the gravitational force between them

"F=\frac{Gmm'}{d^2}"

Putting all values

"F=\frac{6.67\times10^{-11}\times1000\times50}{(0.3)^2}\\=3.7mN"


Now reading or change in mass is given by

"\Delta m=\frac{F}{g}\\=0.38g"